OpenStudy (anonymous):

All of the following equations have the same solution except _____. 3m + 2 = -10 2 - 3m = -10 -10 = -3m + 2 -3m + 2 = -10

OpenStudy (anonymous):

Find the value of m for all equations. all will have same value except one

OpenStudy (anonymous):

so that would be 10?

OpenStudy (anonymous):

or -10?

OpenStudy (anonymous):

\[3m + 2 = -10\] \[3m = -10 -2\] \[m = -12/3\] similarly find the value of m for all equations and then check which equation is odd one out